# Homophily Enhanced Graph Domain Adaptation

### v2

@@ 1 Introduction
In the era of massive graph data collection, graphs often integrate both structural topology and node attributes, providing rich contexts for numerous real-world applications. However, they are often constrained by label scarcity, as annotating structured data remains challenging and costly  (Xu et al., 2022, 2024; Zeng et al., 2024, 2023). To address this challenge, Graph Domain Adaptation (GDA) has emerged as an effective paradigm to transfer knowledge from labeled source graphs to unlabeled target graphs  (Chen et al., 2019; Shi et al., 2024). Conventional GDA methods primarily focus on aligning graph structures across domains by leveraging deep domain adaptation techniques  (Shen et al., 2020b; Wu et al., 2020). While these approaches have achieved promising results in GDA  (Yan & Wang, 2020; Shen et al., 2023), their efforts are centered on mitigating the discrepancy in structural and attribute distributions.
@@ 2.1 Heterophilic Graph Learning
Heterophilic structure is prevalent in practice, from personal relationships in daily life to chemical and molecular scientific study (Fang et al., 2022; He et al., 2025). Developing powerful heterophilic GNN models is a hot research topic.  (Lim et al., 2021; Yang et al., 2024a, 2023b; Zhuo et al., 2024c) provide general benchmarks for heterophilic graph learning. In addition, many methods have been proposed to revise GNNs for heterophilic graphs.  (Yang et al., 2021) specifies propagation weight for each attribute to make GNNs fit heterophilic graphs and  (Li et al., 2022a) explores the underlying homophilic information by capturing the global correlation of nodes.  (Zhu et al., 2020a) enlarges receptive field via exploring high-order structure.  (Chien et al., 2021) adaptively combines the representation of each layer and  (Chen et al., 2020) integrates embeddings from different depths with residual operation. Recent studies  (Mao et al., 2024a; Luan et al., 2022; Huang et al., 2023, 2024; Zhuo et al., 2025, 2024b) reveal that graph homophilc and heterophilic patterns impact graph clustering performance between the test and training sets. However, the above methods failed to explore the role of homophily in GDA and minimize their discrepancy in many aspects (Pu et al., 2025; Xu et al., 2025).
@@ 2.2 Graph Domain Adaptation
Recent Domain Adaptation works have differences from GDA methods(Li et al., 2024a, 2025b, b, 2025a). Identifying the differences between the target and source graphs in GDA is crucial. For graph-structured data, several studies have explored cross-graph knowledge transfer using graph domain adaptation (GDA) methods (Shen & Chung, 2019; Dai et al., 2022; Shi et al., 2024). Some graph inforamtion alignment-based methods (Shen et al., 2020a, b; Yan & Wang, 2020; Shen et al., 2023; Yang et al., 2023a) adapt graph source node label information by integrating global and local structures from both nodes and their neighbors. UDAGCN (Wu et al., 2020) introduces a dual graph convolutional network that captures both local and global knowledge, adapting it through adversarial training. Furthermore, ASN and GraphAE (Zhang et al., 2021; Guo et al., 2022) consider extracting and aligning graph specific information like node degree and edge shift, enabling the extraction of shared features across networks. SOGA (Mao et al., 2024b) is the first to incorporate discriminability by promoting structural consistency between target nodes of the same class, specifically for source-free domain adaptation (SFDA) on graphs. SpecReg (You et al., 2022) applies an optimal transport-based GDA bound and demonstrates that revising the Lipschitz constant of GNNs can enhance performance through spectral smoothness and maximum frequency response. JHGDA (Shi et al., 2023) tackles hierarchical graph structure shifts by aggregating domain discrepancies across all hierarchy levels to provide a comprehensive discrepancy measure. ALEX (Yuan et al., 2023) creates a label-shift-enhanced augmented graph view using a low-rank adjacency matrix obtained through singular value decomposition, guided by a contrasting loss function. SGDA (Qiao et al., 2023) incorporates trainable perturbations (adaptive shift parameters) into embeddings via adversarial learning, enhancing source graphs and minimizing marginal shifts. PA (Liu et al., 2024c) mitigates structural and label shifts by recalibrating edge weights to adjust the influence among neighboring nodes, addressing conditional structure shifts effectively.
@@ Appendix A Proof
Proposition 1 (Bound for $D^{\gamma}_{S,T}(P;\lambda)$, Adaptation from   (Ma et al., 2021; Mao et al., 2024a; anonymous2025on)).
@@
From the cited works (Ma et al., 2021; Mao et al., 2024a; anonymous2025on), it follows that the expected loss discrepancy measure $D^{\gamma/2}_{S,T}(P;\lambda)$ can be bounded by:

### v3

@@ 1 Introduction
In the era of massive graph data collection, graphs often integrate both structural topology and node attributes, providing rich contexts for numerous real-world applications. However, they are often constrained by label scarcity, as annotating structured data remains challenging and costly  (Xu et al., 2022, 2024b; Zeng et al., 2024, 2023). To address this challenge, Graph Domain Adaptation (GDA) has emerged as an effective paradigm to transfer knowledge from labeled source graphs to unlabeled target graphs  (Chen et al., 2019; Shi et al., 2024). Conventional GDA methods primarily focus on aligning graph structures across domains by leveraging deep domain adaptation techniques  (Shen et al., 2020b; Wu et al., 2020). While these approaches have achieved promising results in GDA  (Yan & Wang, 2020; Shen et al., 2023), their efforts are centered on mitigating the discrepancy in structural and attribute distributions.
@@ 2.1 Heterophilic Graph Learning
Heterophilic structure is prevalent in practice, from personal relationships in daily life to chemical and molecular scientific study (Fang et al., 2022; He et al., 2025). Developing powerful heterophilic GNN models is a hot research topic.  (Lim et al., 2021; Qian et al., 2024; Yang et al., 2024a, 2023b; Zhuo et al., 2024c; Li et al., 2025b) provide general benchmarks for heterophilic graph learning. In addition, many methods have been proposed to revise GNNs for heterophilic graphs.  (Yang et al., 2021) specifies propagation weight for each attribute to make GNNs fit heterophilic graphs and  (Li et al., 2022a) explores the underlying homophilic information by capturing the global correlation of nodes.  (Zhu et al., 2020a) enlarges receptive field via exploring high-order structure.  (Chien et al., 2021) adaptively combines the representation of each layer and  (Chen et al., 2020) integrates embeddings from different depths with residual operation. Recent studies  (Mao et al., 2024a; Liu et al., 2025; Li et al., 2024a, 2025a; Luan et al., 2022; Huang et al., 2023, 2024; Zhuo et al., 2025, 2024b) reveal that graph homophilc and heterophilic patterns impact graph clustering performance between the test and training sets. However, the above methods failed to explore the role of homophily in GDA and minimize their discrepancy in many aspects (Kang et al., 2024; Pu et al., 2025; Xu et al., 2025; Xie et al., 2024, 2025).
@@ 2.2 Graph Domain Adaptation
Recent Domain Adaptation works have differences from GDA methods  (Li et al., 2024b; Chen et al., 2024; Li et al., 2025c, 2024c, b; Zeng et al., 2025; Xu et al., 2024a). Identifying the differences between the target and source graphs in GDA is crucial. For graph-structured data, several studies have explored cross-graph knowledge transfer using graph domain adaptation (GDA) methods (Shen & Chung, 2019; Dai et al., 2022; Shi et al., 2024). Some graph information alignment-based methods (Shen et al., 2020a, b; Yan & Wang, 2020; Shen et al., 2023; Yang et al., 2023a) adapt graph source node label information by integrating global and local structures from both nodes and their neighbors. UDAGCN (Wu et al., 2020) introduces a dual graph convolutional network that captures both local and global knowledge, adapting it through adversarial training. Furthermore, ASN and GraphAE (Zhang et al., 2021; Guo et al., 2022) consider extracting and aligning graph specific information like node degree and edge shift, enabling the extraction of shared features across networks. SOGA (Mao et al., 2024b) is the first to incorporate discriminability by promoting structural consistency between target nodes of the same class, specifically for source-free domain adaptation (SFDA) on graphs. SpecReg (You et al., 2022) applies an optimal transport-based GDA bound and demonstrates that revising the Lipschitz constant of GNNs can enhance performance through spectral smoothness and maximum frequency response. JHGDA (Shi et al., 2023) tackles hierarchical graph structure shifts by aggregating domain discrepancies across all hierarchy levels to provide a comprehensive discrepancy measure. ALEX (Yuan et al., 2023) creates a label-shift-enhanced augmented graph view using a low-rank adjacency matrix obtained through singular value decomposition, guided by a contrasting loss function. SGDA (Qiao et al., 2023) incorporates trainable perturbations (adaptive shift parameters) into embeddings via adversarial learning, enhancing source graphs and minimizing marginal shifts. PA (Liu et al., 2024c) mitigates structural and label shifts by recalibrating edge weights to adjust the influence among neighboring nodes, addressing conditional structure shifts effectively.
@@ Appendix A Proof
Proposition 1 (Bound for $D^{\gamma}_{S,T}(P;\lambda)$, Adaptation from   (Ma et al., 2021; Mao et al., 2024a; Fang et al., 2025)).
@@
From the cited works (Ma et al., 2021; Mao et al., 2024a; Fang et al., 2025), it follows that the expected loss discrepancy measure $D^{\gamma/2}_{S,T}(P;\lambda)$ can be bounded by: